### v5

# Sinaĭ excursions: An analogue of Sparre Andersen’s formula for the area process of a random walk

Abstract.
Sinaĭ initiated the study of random walks with persistently positive area processes, motivated by shock waves in solutions to the inviscid Burgers’ equation. We find the precise asymptotic probability that the area process of a random walk bridge is an excursion. A key ingredient is an analogue of Sparre Andersen’s classical formula. The asymptotics are related to von Sterneck’s subset counting formulas from additive number theory. Our results sharpen bounds by Aurzada, Dereich and Lifshits and respond to a question of Caravenna and Deuschel, which arose in their study of the wetting model. In this context, Sinaĭ excursions are a class of random polymer chains exhibiting entropic repulsion.
Key words and phrases:
entropic repulsion; infinite divisibility; majorization; persistence probability; random polymer; random walk; renewal sequence; wetting model
2010 Mathematics Subject Classification:
05A15; 05A17; 05C20; 60E07; 60G50; 60K05; 82B41; 82D60
1. Introduction
1.1. Persistence probabilities
There is a rich history of calculating persistence probabilities in mathematics, wherein we ask for a random process to continue to satisfy some property of interest. For instance, Bertrand’s ballot theorem [1] from 1887 can be viewed as an early example. No less famous is Sparre Andersen’s [28] formula from 1954, concerning the probability that a simple random walk remains positive.
Applications of persistence in statistical physics began in the late 1980s, as it relates to, e.g., fluctuating interfaces and sticky particle systems. We refer to the surveys [23, 13, 10, 3] for a detailed overview of existing results and applications.
In this work, we introduce an analogue of Sparre Andersen’s formula for the area process of a random walk; see (2) and Theorem 1.2 below.
Persistence probabilities for such area processes were first studied by Sinaĭ [26], in relation to the inviscid Burgers’ equation. This equation models a turbulent fluid which gives rise to shock waves. Such discontinuities are overcome mathematically via control surfaces, i.e., interfaces through which material flows from one side of a discontinuity to the other. When the system is started with self-similar, Brownian data, the exponent $1/4$ in (1) below is related to the fact that the set of initial positions of particles not yet “shocked” by time $t=1$ has Hausdorff dimension $1/2$; see [3, 27, 8, 21].
Times when the random walk and its area process revisit 0 are renewal times for such persistence problems. We call the trajectories between such times Sinaĭ excursions; see Section 1.3 below. These excursions play a key role in the work of Caravenna and Deuschel [11] on the wetting model. In this context, Sinaĭ excursions are related to random polymer chains exhibiting a phenomenon referred to as entropic repulsion. More specifically, the area process of a Sinaĭ excursion is used to model the interface that forms between a gas pressed diffusively against a surface by a liquid; see [11, 3, 9, 29, 20] for more details.
Our main result Theorem 1.1 (see also Corollary 5.1) identifies the precise asymptototic probability that a random walk is a Sinaĭ excursion. This sharpens the original bounds (up to polylogarithmic factors) by Caravenna and Deuschel [11] and the subsequent improvement (up to constant factors) by Aurzada, Dereich and Lifshits[2]. Our proof utilizes our area analogue of Sparre Andersen’s formula mentioned above, old formulas from additive number theory proved in the early 1900s by von Sterneck [4], and the Tauberian theorems related to the Lévy–Khintchine formula proved by Hawkes and Jenkins [22].
1.2. Sinaĭ walks
Let $(S_{k}:k\geqslant 0)$ be a simple symmetric random walk on the integers $\mathbb{Z}$ started at $S_{0}=0$. We let $A_{k}=\sum_{i=1}^{k}S_{i}$ denote its cumulative area after $k$ steps. Sinaĭ [26] proved that
${\bf P}(A_{1},\ldots,A_{n}\geqslant 0)=\Theta(n^{-1/4}).$ (1)
If $A_{1},\ldots,A_{n}\geqslant 0$ holds, we call $(S_{0},S_{1},\ldots,S_{n})$ a Sinaĭ walk.
As discussed by Aurzada and Simon [3, Section 3], in their survey on persistence probabilities, Sinaĭ’s proof is based on the sequence of times $0=\tau_{0},\tau_{1}\ldots$ that the walk visits 0. This gives rise to another random walk, whose increments are the signed areas accumulated between these times. A key ingredient is Sparre Andersen’s [28] classical result that, for $|x|\leqslant 1$,
$\sum_{n=0}^{\infty}{\bf P}(T_{0}>n)x^{n}=\exp\left(\sum_{k=1}^{\infty}{\bf P}(S_{k}\leqslant 0)\frac{x^{k}}{k}\right),$ (2)
where $T_{0}=\inf\{t:S_{t}>0\}$ is the first time that the walk is positive.
The utility of (2) lies in the fact that the probabilities ${\bf P}(S_{k}\leqslant 0)$ are simpler than ${\bf P}(T_{0}>n)$. A Tauberian theorem implies ${\bf P}(T_{0}>\tau_{n})\sim cn^{-1/2}$. Finally, (1) is derived using that $n^{-2}\tau_{n}$ converges to a stable random variable.
We note that Vysotsky [30] sharpened and generalized Sinaĭ’s persistence probability (1), showing that
${\bf P}(A_{1},\ldots,A_{n}\geqslant 0)\sim Cn^{-1/4},$
for a wide class of random walks. The constant $C$, however, is expressed in terms of a rather complicated integral (see equation (37) therein). Perhaps our current arguments can help with finding a more explicit, combinatorial description of $C$, at least in some cases.
More specifically, in [30] it is assumed that the increments of the random walk are $\alpha$-stable, for some $1<\alpha\leqslant 2$, and either right-exponential or right-continuous (see [30] for the precise conditions). Subsequently, Dembo, Ding and Gao [19] extended these results to the case that the increments have finite, positive second moment. See also Gao, Liu and Yang [19] for the case of Gaussian increments.
Finally, we note that continuous analogues of such persistence problems have been studied very recently by Bär, Duraj and Wachtel [19], giving rise to a class of random processes called Kolmogorov diffusions.
Figure 1. The $16$ Sinaĭ excursions of length $12$, slightly staggered. Such excursions have total area 00 and non-negative partial areas. The standard Sinaĭ excursion in bold oscillates between ±$\pm 1$.
1.3. Sinaĭ excursions
We call $(S_{0},S_{1},\ldots,S_{4n})$ a Sinaĭ excursion if it is a Sinaĭ walk and $S_{4n}=A_{4n}=0$. Such renewal times are only possible at multiples of $4$. We note that Sinaĭ excursions are the discrete analogue of the positive Kolmogorov excursions (from zero and back) studied in [19].
Aurzada, Dereich and Lifshits [2] showed that
$p_{n}={\bf P}(A_{1},\ldots,A_{4n}\geqslant 0\mid S_{4n}=A_{4n}=0)=\Theta(n^{-1/2}),$ (3)
verifying a conjecture in Caravenna and Deuschel [11].
Our main result identifies the precise asymptotics.
Theorem 1.1.
As $n\to\infty$,
$n^{1/2}p_{n}\to\frac{1}{2}\sqrt{\frac{\pi}{6}}\exp\left(\sum_{k=1}^{\infty}\frac{\Xi_{k}}{k2^{4k}}\right),$
where
$\Xi_{k}=\frac{1}{4n}\sum_{d|2k}{4k/d-1\choose 2k/d}\phi(d),$ (4)
and $\phi$ is Euler’s totient function.
Our methods also lead to the precise asymptotics of
${\bf P}(A_{1},\ldots,A_{2n}\geqslant 0\mid S_{2n}=0),$ (5)
corresponding to a class of Sinaĭ meanders. See Corollary 5.1 below.
As we will discuss in Section 2 below, $\Xi_{n}$ is the number of subsets of $\{1,2,\ldots,4n-1\}$ of size $2n$ that sum to $3n$ mod $4n$. This formula is one instance in a family of general modular subset counting formulas proved by von Sterneck in the early 1900s.
To give a first hint about the connection between $p_{n}$ and $\Xi_{n}$, consider the $2n$ times $t_{1}<\cdots<t_{2n}$ before down steps in a walk $(S_{0},S_{1},\ldots,S_{4n})$ when $S_{t+1}-S_{t}=-1$. By Lemma 3.1 below, if $A_{4n}=0$ then $\sum_{j=1}^{2n}t_{j}=n(4n-1)$. Thus, $\{t_{1},\ldots,t_{2n}\}$ is a subset of $\{1,2,\ldots,4n-1\}$ of size $2n$ that sums to $3n$ mod $4n$. More to the point, as we will see, $\Xi_{n}$ is related to the number of bridges that can be turned into a Sinaĭ excursion, by cyclically shifting their increments.
1.4. A Sparre Andersen analogue
To prove Theorem 1.1, we will first establish the following analogue of Sparre Andersen’s formula (2) for the probabilities
$\varphi_{n}={\bf P}(A_{1},\ldots,A_{4n}\geqslant 0,\>A_{4n}=S_{4n}=0).$ (6)
By the local limit in [2, Proposition 2.1], it follows that
$n^{2}{\bf P}(A_{4n}=S_{4n}=0)\to\frac{\sqrt{3}}{4\pi}.$
Therefore, to prove Theorem 1.1, it suffices to show that
$n^{5/2}\varphi_{n}\to\frac{e^{\lambda}}{8\sqrt{2\pi}},$ (7)
where
$\lambda=\sum_{k=1}^{\infty}\frac{\Xi_{k}}{k2^{4k}}.$ (8)
For convenience, we put $\xi_{n}=\Xi_{n}/2^{4n}$ and let $\Phi_{n}=2^{4n}\varphi_{n}$ denote the number of Sinaĭ excursions of length $4n$.
Theorem 1.2.
For $|x|\leqslant 1$,
$\sum_{n=0}^{\infty}\varphi_{n}x^{n}=\exp\left(\sum_{k=1}^{\infty}\xi_{k}\frac{x^{k}}{k}\right).$ (9)
As with (2), the usefulness of (9) is that it allows for an indirect analysis of the probabilities of interest $\varphi_{n}$. The criteria for Sinaĭ excursions imposes conditions at all times along the trajectory. On the other hand, $\xi_{n}$ has a much simpler description, by von Sterneck’s formulas (Lemma 2.1 below).
1.5. Transferring asymptotics
To transfer asymptotic information from $\xi_{n}$ to $\varphi_{n}$, we will use a connection with Lévy processes $(L_{t},t\geqslant 0)$. We recall that a probability measure $\pi$ is infinitely divisible if for all $m\geqslant 1$ there are independent and identically distributed $X_{1},\ldots,X_{m}$ for which $X_{1}+\cdots+X_{m}\sim\pi$. Lévy processes have independent, stationary increments, so the distribution of $L_{t}$ is infinitely divisible, at any given time $t>0$.
The Lévy–Khintchine formula relates an infinitely divisible $\pi$ to a certain Lévy measure $\nu$, which controls the jumps in the associated Lévy process $(L_{t},t\geqslant 0)$ such that $L_{1}\sim\pi$. In the case that $\pi=(p_{n},n\geqslant 0)$ is supported on the non-negative integers, we have that
$\sum_{n=0}^{\infty}p_{n}x^{n}=\exp\left(\sum_{k=1}^{\infty}(1-x^{k})\nu_{k},\right)$ (10)
where $(\nu_{k},k\geqslant 1)$ has finite total Lévy measure $\lambda=\sum_{k=1}^{\infty}\nu_{k}$. Furthermore, $\nu_{k}$ is the expected number of jumps of size $k$ by time $t=1$.
Hence, by Theorem 1.2, it follows that $p_{n}=e^{-\lambda}\varphi_{n}$ is infinitely divisible, with corresponding Lévy measure $\nu_{n}=\xi_{n}/n$. The constant $\lambda$ in (8), which appears in Theorem 1.1, is the total Lévy measure. To complete the proof of Theorem 1.1, given Theorem 1.2, we will use the following result by Embrechts and Hawkes [16], which shows that $p_{n}\sim\nu_{n}$, when $\nu_{n}$ is sufficiently regular.
A probability distribution $(q_{n},n\geqslant 0)$ is sub-exponential if $q_{n}/q_{n+1}\to 1$ and its convolution
$q_{n}^{*}=\sum_{k=0}^{n}q_{k}q_{n-k}.$
satisfies that $q_{n}^{*}/q_{n}\to 2$. In [16] it is proved that if $p_{n}$ and $\nu_{n}$ are related by (10), then $\nu_{n}/\lambda$ is sub-exponential if and only if $p_{n}\sim\nu_{n}$ and $\nu_{n}/\nu_{n+1}\to 1$. Intuitively, this follows by the “one big jump principle.” A large value of $L_{1}$ is likely due to one big jump of essentially this value by time $t=1$. As a special case, Hawkes and Jenkins [22] showed that $p_{n}\sim\nu_{n}$ if $\nu_{n}$ is regularly varying with index $\gamma<-1$.
Let us note that the essential feature of $\exp(z)$ in (10) is that it is analytic. Indeed, the results in [16] are based on the work of Chover, Ney and Wainger [12] on analytic transformations of probability measures, and so extend to other analytic $f(z)$; see Embrechts and Omey [17].
In the present case, by Stirling’s approximation,
$n^{5/2}\nu_{n}=n^{3/2}\xi_{n}=\frac{n^{3/2}}{2^{4n}}\Xi_{n}\to\frac{1}{8\sqrt{2\pi}},$
since the term $d=1$ dominates in (4). As such, $\nu_{n}$ is regularly varying with index $\gamma=-5/2$, and it follows that
$n^{5/2}\varphi_{n}=e^{\lambda}n^{5/2}p_{n}\sim e^{\lambda}n^{5/2}\nu_{n}\to\frac{e^{\lambda}}{8\sqrt{2\pi}},$
yielding (7). As discussed, Theorem 1.1 follows.
1.6. Outline
We have shown how Theorem 1.1 follows from Theorem 1.2 and the asymptotics of $\Xi_{n}$. The remainder of the article is devoted to the proof of Theorem 1.2.
2. von Sterneck’s formulas
In the early 1900s, von Sterneck (see, e.g., [4, 24]) found the number $\Lambda_{k}(n,s)$ of multi-sets $\{m_{1},\ldots,m_{k}\}$ of $\{0,1,\ldots,n-1\}$ of size $k$ that sum to $\sum_{i=1}^{k}m_{i}\equiv s$ mod $n$.
Lemma 2.1 (von Sterneck).
For all $n\geqslant 1$, we have that
$\Lambda_{k}(n,s)=\frac{1}{n}\sum_{d|k,n}{(n+k)/d-1\choose k/d}\frac{\mu(d/\gcd(d,s))\phi(d)}{\phi(d/\gcd(d,s))},$
where $\mu$ is the Möbius function, $\phi$ is the Euler totient function, and $\gcd(d,s)$ is the greatest common divisor of $d$ and $s$.
In particular,
$\Lambda_{2n}(2n,0)=\frac{1}{n}\sum_{d|2n}{4n/d-1\choose 2n/d}\phi(d).$
Therefore, to justify (4) above, we prove the following claim.
Lemma 2.2.
For all $n\geqslant 1$, we have that
$2\Xi_{n}=\Lambda_{2n}(2n,0).$
Proof.
First, we note that, for integers $11\leqslant a_{1}<\cdots<a_{2n}\leqslant 4n-$ 1, we have $\sum_{i=1}^{2n}a_{i}\equiv n$ mod $4n$ if and only if $\sum_{i=1}^{2n}(4n-a_{i})\equiv 3n$ mod $4n$. Hence, there is the same number of subsets of $\{1,2,\ldots,4n-1\}$ of size $2n$ that sum to $n$ mod $4n$ as there are that sum to $3n$ mod $4n$.
Next, we claim that, for integers $10\leqslant m_{1}\leqslant\cdots\leqslant m_{2n}\leqslant 2n-$ 1, we have that $\sum_{i=1}^{2n}m_{i}$ is equal to 00 mod $2n$ if and only if $\sum_{i=1}^{2n}(m_{i}+i)$ is equal to $n$ or $3n$ mod $4n$. Indeed, to see this, simply note that $\sum_{i=1}^{2n}i=n(2n+1)$ is equal to $n$ or $3n$ mod $4n$ (depending on the parity of $n$). This implies that the sub-multisets of $\{1,\dots,2n-1\}$ of size $2n$ that sum to 00 mod $2n$ are in bijection with the subsets of $\{1,\dots,4n-1\}$ of size $2n$ that sum to $n$ or $3n$ mod $n$, and the statement follows. ∎
3. Times before down steps
Consider a bridge ${\mathcal{B}}=(B_{0},B_{1},\ldots,B_{2n})$ of length $2n$. That is, $B_{0}=B_{2n}=0$ and all increments $\Delta B_{k}=B_{k+1}-B_{k}=\pm 1$, for $10\leqslant k\leqslant 2n-$ 1. Let
${\bf t}({\mathcal{B}})=(t_{1},\ldots,t_{n})$
denote the sequences of times before down steps, that is, times $10\leqslant t_{1}<\dots<t_{n}\leqslant 2n-$ 1 such that $\Delta B_{t}=-1$.
Lemma 3.1.
Let ${\mathcal{B}}=(B_{0},B_{1},\ldots,B_{4n})$ be a bridge with times ${\bf t}({\mathcal{B}})=(t_{1},\ldots,t_{2n})$ before down steps. Then its total area
$A_{4n}=\sum_{k=1}^{4n}B_{k}=-2n(4n-1)+2\sum_{i=1}^{2n}t_{i}.$ (11)
Proof.
To see this, note that
$\sum_{k=1}^{4n}B_{k}=\sum_{k=0}^{4n-1}(4n-k)\Delta B_{k}=\sum_{k=0}^{4n-1}(4n-k)-2\sum_{j=1}^{2n}(4n-t_{j}),$
where the last step follows from $\Delta B_{k}=1-2\mathbf{1}{\{k\in\{t_{1},\dots,t_{2n}\}\}}$. This simplifies to $-2n(4n-1)+2\sum_{i=1}^{2n}t_{i}$, as claimed. ∎
We let
$\mathcal{S}=(0,1,0,-1,0\ldots,0,1,0,-1,0)$
denote the standard Sinaĭ excursion of length $4n$; see Figure 1. Note that $\mathcal{S}$ is a “sawtooth” bridge, oscillating between ±$\pm 1$, with
${\bf t}(\mathcal{S})=(1,2,5,6,\ldots,4n-3,4n-2).$ (12)
Note that ${\bf t}(\mathcal{S})$ sums to $n(4n-1)$. Therefore, if ${\mathcal{B}}$ is a Sinaĭ excursion then, by Lemma 3.1, ${\bf t}({\mathcal{B}})$ and ${\bf t}(\mathcal{S})$ have the same sum.
In fact, it can be shown that ${\mathcal{B}}$ is a Sinaĭ excursion if and only if ${\bf t}({\mathcal{B}})$ is majorized by ${\bf t}(\mathcal{S})$. (For weakly increasing ${\bf x},{\bf y}\in\mathbb{R}^{n}$, ${\bf x}$ is majorized by ${\bf y}$ if all partial sums $\sum_{i=1}^{k}x_{i}\geqslant\sum_{i=1}^{k}y_{i}$, with equality when $k=n$.) Intuitively, $\mathcal{S}$ takes its down steps as soon as possible, maintaining a cumulative area $\geqslant 0$ as close to 0 as possible. We will not require this fact, and omit the details.
4. Sparre Andersen for Sinaĭ excursions
In this section, we prove Theorem 1.2, which states that
$\sum_{n=0}^{\infty}\varphi_{n}x^{n}=\exp\left(\sum_{k=1}^{\infty}\xi_{k}\frac{x^{k}}{k}\right).$
By differentiating and comparing coefficients, it can be seen that this is equivalent to
$n\varphi_{n}=\sum_{k=1}^{n}\xi_{k}\varphi_{n-k}.$
Therefore, multiplying by $2^{4n}$, to prove Theorem 1.2 it suffices to show
$n\Phi_{n}=\sum_{k=1}^{n}\Xi_{k}\Phi_{n-k}.$ (13)
For this, we use the observation that $\Phi_{n}$ is a renewal sequence. A renewal sequence enumerates structures that can be decomposed into a series of irreducible parts. Formally, $A_{n}$ is a renewal sequence if its generating function $A(x)=\sum_{n=0}^{\infty}A_{n}x^{n}$ can be expressed as
$A(x)=\sum_{m=0}^{\infty}[A^{(1)}(x)]^{m}=\frac{1}{1-A^{(1)}(x)}$
where $A^{(1)}(x)=\sum_{n=1}^{\infty}A_{n}^{(1)}x^{n}$ is the generating function for the number $A_{n}^{(1)}$ of irreducible structures of length $n$. See, e.g., Feller [18] for details.
We see that $\Phi_{n}$ is a renewal sequence because any Sinaĭ excursion can be decomposed into a series of irreducible Sinaĭ excursions whose area process only takes the value 00 at the first and the last step. More specifically, let $\Phi_{n}^{(1)}$ be the number of walks $(S_{0},S_{1},\ldots,S_{4n})$ for which $S_{0}=A_{0}=0$, $S_{4n}=A_{4n}=0$ and $A_{1},\ldots,A_{4n-1}>0$. Then, the generating function $\Phi(x)=\sum_{n}\Phi_{n}x^{n}$ of $\Phi_{n}$ can be expressed as
$\Phi(x)=\frac{1}{1-\Phi^{(1)}(x)},$
where $\Phi^{(1)}(x)=\sum_{n}\Phi_{n}^{(1)}x^{n}$ is the generating function for the number $\Phi^{(1)}_{n}$ of irreducible Sinaĭ excursions.
Then, the key to proving (13) is the following lemma, proved by Bassan and the authors in [7].
Lemma 4.1 (Lemma 2 (1) in [7]).
Suppose that $1=A_{0},A_{1},\ldots$ is a renewal sequence, where $A_{n}$ counts the number of objects in a class $\mathcal{A}_{n}$ of objects of size $n$. Then
$nA_{n}=\sum_{k=1}^{n}A_{k}^{\prime}A_{n-k},$ (14)
where $A_{n}^{\prime}$ is the number of pairs $(X,s)$, where $X\in\mathcal{A}_{n}$ and $0\leqslant s<\ell$ is an integer, where $\ell$ is the size of the first irreducible part of $X$.
Its proof is simple, so we include it here.
Proof.
Note that $nA_{n}$ enumerates the pairs $(X,i)$, where $X\in\mathcal{A}_{n}$ and $0\leqslant i<n$ is an integer. We call such pairs marked objects of size $n$. For a marked object $(X,i)$ of size $n$, consider the sub-object consisting of the irreducible part containing $i$ and all subsequent parts. This sub-object is of size $k$, for some $1\leqslant k\leqslant n$, and has a mark somewhere in its first irreducible part. All other previous parts form an (unmarked) object of size $n-k$. This implies that there are $\sum_{k=1}^{n}A_{k}^{\prime}A_{n-k}$ many such pairs $(X,i)$. See Figure 2. ∎
Figure 2. Marked objects of size $n$ ($nA_{n}$ many) can be split into an object of size $n-k$ ($A_{n-k}$ many) and an object of size $k$ with a mark in its first irreducible part ($A_{k}^{\prime}$ many). Vertical lines delimit irreducible parts. Recurrence (14) follows, summing over $k$.
Lemma 4.1 and the observation that $\Phi_{n}$ is a renewal sequence imply the following lemma (recalling that renewal times when $S_{4k}=A_{4k}=0$ are only possible at multiples of $4$).
Lemma 4.2.
Let $\Phi^{\prime}_{n}$ denote the number of pairs $({\mathcal{B}},s)$, where ${\mathcal{B}}$ is a Sinaĭ excursion of length $4n$ and $0\leqslant s<\ell$ is an integer, where $\ell$ the length of the first irreducible Sinaĭ excursion of ${\mathcal{B}}$. Then,
$n\Phi_{n}=\sum_{k=1}^{n}\Phi_{k}^{\prime}\Phi_{n-k}.$
Finally, to prove (13) and thereby complete the proof of Theorem 1.2, we show the following.
Lemma 4.3.
For all $n\geqslant 1$, we have that $\Phi_{n}^{\prime}=\Xi_{n}$.
Proof.
In fact, it will be easier to show that $2\Phi^{\prime}_{n}=2\Xi_{n}$. By Lemma 2.2, this is the number of subsets of $\{1,2,\ldots,4n-1\}$ of size $2n$ that sum to $n$ or $3n$ mod $4n$. To do this, we will find a bijection $\Upsilon$
  • from the set of all pairs $({\mathcal{B}},j)$, where ${\mathcal{B}}$ is a Sinaĭ excursion of length $4n$, with first positive Sinaĭ excursion of length $\ell=4k$, and $1\leqslant j\leqslant 4k$ is an integer,
  • to the set of all of subsets $T$ of $\{1,2,\ldots,4n-1\}$ of size $2n$ that sum to $n$ or $3n$ mod $4n$.
To describe $\Upsilon$, consider ${\mathcal{B}}$ as above. Let
$10=i_{1}<\cdots<i_{2k}=4k-$ 1
be the times before up steps in the first positive Sinaĭ excursion of ${\mathcal{B}}$. Let ${\mathcal{B}}^{(j)}$ be the bridge obtained from ${\mathcal{B}}$ by cyclically shifting ${\mathcal{B}}$ to the left by $i_{j}$. In other words, the $k$th increment of ${\mathcal{B}}^{(j)}$ is the $(k+i_{j})$th (understood mod $4n$) increment of ${\mathcal{B}}$. In particular, ${\mathcal{B}}^{(1)}={\mathcal{B}}$. We let $\Upsilon({\mathcal{B}},j)$ to be the set of times before down steps in ${\mathcal{B}}^{(j)}$. Lemma 3.1 implies that, because ${\mathcal{B}}$ is a Sinaĭ excursion, its times before down steps sum to $n(4n-1)\equiv 3n$ mod $4n$. We shift all of these $2n$ times by the same amount to obtain the times before down steps in ${\mathcal{B}}^{(j)}$, so these sum to $n$ or $3n$ mod $4n$. See Figure 3 for an example.
Finally, let us describe $\Upsilon^{-1}$. Let $T$ be as above. Consider the bridge ${\mathcal{X}}$ with times before down steps at times $t\in T$. Then, by Lemma 3.1, the total area $A$ of ${\mathcal{X}}$ is equal to 00 mod $4n$. If we translate the $x$-axis by some $\delta\in\mathbb{Z}$ the area of ${\mathcal{X}}$, with respect to this new axis, is $A^{\prime}=A-4\delta n$. Select the unique $\delta$ that sets $A^{\prime}=0$. To find $\Upsilon^{-1}(T)$, we choose the rightmost point before an up step in ${\mathcal{X}}$ along this new axis for which the corresponding cyclic shift forms a Sinaĭ excursion (with respect to this new axis). Since the total area is 0, such a point exists by Raney’s lemma [25]. See Figure 4 for an example. If this point occurs at time $m$ we set ${\mathcal{B}}_{i}={\mathcal{X}}_{i+m}$ (with indices modulo $4n$) and we let $j$ be the index such that the $j$th up step in ${\mathcal{B}}$ occurs at time $4n-m$. ∎
Figure 3. 1st row: Times before down steps 1256, 1346 and 2345 are solid dots in the $\Phi_{2}=3$ Sinaĭ excursions of length $8$. Times before up steps 03, 0257 and 0167 in their first positive excursions are open dots. The bijection $\Upsilon$ gives the $102\Xi_{2}=$ 10 subsets of $\{1,2,\ldots,7\}$ of size $4$ that sum to $2$ or $6$ mod $8$ as follows. 1st column: If we shift the first Sinaĭ excursion by starting at its 1st or 2nd open dot, we obtain bridges with times before down steps (solid dots) at times 1256 and 2367. 2nd column: If we shift the 2nd Sinaĭ excursion by starting at its 1st, 2nd, 3rd or 4th open dot, we obtain bridges with times before down steps at 1346, 1247, 1467 and 2457. 3rd column: If we shift the 3nd Sinaĭ excursion by starting at its 1st, 2nd, 3rd or 4th open dot, we obtain bridges with times before down steps at 2345, 1234, 4567 and 3456.
Figure 4. To invert the bijection $\Upsilon$, we select the horizontal line (dotted) that “cuts” the area in half, and use Raney’s lemma to find the rightmost point (open dot) before an up step that starts a Sinaĭ excursion, with respect to this line.
5. A related application
Finally, let us conclude with another, related application of our current techniques.
As a consequence of considerably more technical arguments than those in the current article, we [15, Corollary 4] recently proved that
$n^{1/2}p_{n}\to\frac{1}{2}\sqrt{\frac{\pi}{6}}\frac{1}{1-{\bf P}(A_{\tau}=0)},$ (15)
where $\tau=\inf\{t:S_{t}=0,\>A_{t}\leqslant 0\}$. Informally, $\tau$ is the first time that a random walk is “at risk” of not being Sinaĭ (i.e., if $A_{\tau}<0$). Theorem 1.1 above implies that ${\bf P}(A_{\tau}=0)=1-e^{-\lambda}$, where $\lambda$ is as in (8), so that
$n^{1/2}p_{n}\to\frac{1}{2}\sqrt{\frac{\pi}{6}}e^{\lambda}.$
This, in turn, when combined with Proposition 6.3 in [5], yields the following corollary, concerning the probability that a random walk bridge of length $2n$ is a Sinaĭ walk.
Corollary 5.1.
As $n\to\infty$, it holds that
$n^{1/4}{\bf P}(A_{1},\ldots,A_{2n}\geqslant 0\mid S_{2n}=0)\to\frac{e^{\lambda/2}\sqrt{\pi}}{\Gamma(1/4)}.$
6. Acknowledgments
SD acknowledges the financial support of the CogniGron research center and the Ubbo Emmius Funds (University of Groningen). BK was partially supported by a Florence Nightingale Bicentennial Fellowship (Oxford Statistics) and a Senior Demyship (Magdalen College). BK gratefully acknowledges Vasu Tewari for indicating the work of von Sterneck.